Need help please and thank you.
Chris and Hal live 72 miles apart in southeastern Illinois. They decide to bicycle towards each other and meet somewhere in between. Hal's rate of speed is 60% of Chris's. They start out at the same time and meet 5 hours later. Find Hal's rate of speed.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Chris and Hal live 72 miles apart in southeastern Illinois. They decide to bicycle towards each other and meet somewhere in between. Hal's rate of speed is 60% of Chris's. They start out at the same time and meet 5 hours later. Find Hal's rate of speed.
==============
There is no "rate of speed."
There's rate, and there's speed, but no rate of speed.
-----------------
Confirmed by the Dept. of Redundancy Dept.

Answer by ikleyn(52906)

.

Let x be the Chris' rate, in miles per hour.


Then Hal' rate is 0.6x mph.


In 5 hours, Chris traveled 5x miles;  Hal traveled 5*(0.6x) = 3x miles.


The total distance is


    5x + 3x = 72 miles,   or


    8x = 72

     x = 72/8 = 9.


Thus Chris' rate is 9 mph;  Hal's rate os 0.6*9 = 5.4 mph.


ANSWER.  Hal's rate is 5.4 miles per hour.

Solved.

-----------------

See introductory lessons on Travel and Distance problems
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.